OpenStudy (anonymous):

what is the units digit of 3^3887?

Directrix (directrix):

First look at some powers of 3 and observe the units digit.

Directrix (directrix):

Chart source: http://mathcentral.uregina.ca/QQ/database/QQ.09.05/alex1.html

OpenStudy (ybarrap):

3^0=1 3^1 = 3 3^2=9 3^3=27 3^4=81 3^5=243 3^6=729 3^7=2187 3^8=6561 3^9=19683 Pattern of units digit is 1,3,9,7,1,3,9,7,1,3,9,7,... Every exponent with the same Units digit is separated from each other by 4 units in the exponent: 3^0,3^4,3^8 have units of 1 3^1,3^5,3^9 have units of 3 3^2,3^6 have units of 9 3^3,3^7 have units of 7 And so on. Conjecture: If the number is a perfect 4th root then units digit is a 1 If the number is a perfect cube root then units digit is a 7 If the number is a perfect square root then units digit is a 9 Otherwise, units digit is a 3 For example \(\sqrt{729}=27\), so last digit is a 9. For our case 3^3887 is a perfect cube, so units digit must be a 7.

OpenStudy (ybarrap):

Of course, 3^3887 is a perfect cube. We need to work on the conjecture a little more. I am not certain the order we need to take them yet so I'm not certain on the units digit. Maybe you can take something from this, maybe not.

OpenStudy (ybarrap):

New Conjecture: Take the exponent of 3^N If N mod 4 = 0 then Units digit is a 1 If N mod 4 = 1 then Units digit is a 3 If N mod 4 = 2 then Units digit is a 9 If N mod 4 = 3 then Units digit is a 7 In our case N = 3887 So 3887 mod 4 = 3, so Units digit is a 7 (after all).

OpenStudy (anonymous):

7 is correct. Refer to the complete expansion attached. Calculated and formatted in 33 micro seconds by Mathemtica v9.